# High-mobility heavy quasiparticles in a van der Waals antiferromagnetic dense Kondo lattice CeTe3

### v1

@@
[…]
High-mobility heavy quasiparticles in a van der Waals antiferromagnetic dense Kondo lattice CeTe3
Hai Zeng1∗, Yang Zhang1∗, Bingke Ji2111These authors contribute equally to this work, Jiaqiang Cai3, Shuo Zou1, Zhuo Wang1, Chao Dong1, Kangjian Luo1, Yang Yuan1, Kai Wang4, Jinglei Zhang3, Chuanyin Xi3, Junfeng Wang1, Yaomin Dai2222Electronic address: ymdai@nju.edu.cn, Jing Li1333Electronic address: jing_li@hust.edu.cn, and Yongkang Luo1,5444Electronic address: mpzslyk@gmail.com 1Wuhan National High Magnetic Field Center and School of Physics, Huazhong University of Science and Technology, Wuhan 430074, China; 2School of Physics, Nanjing University, Nanjing 210023, China; 3Anhui Key Laboratory of Low-Energy Quantum Materials and Devices, High Magnetic Field Laboratory, Hefei Institutes of Physical Science, Chinese Academy of Sciences, Hefei 230031, China; 4School of Physics and Electronics, Henan University, Kaifeng 475004, China; 5Lead contact
(June 30, 2025)
SUMMARY: Two-dimensional van der Waals (vdW) materials exhibit high carrier mobility and tunability, making them suitable for low-power, high-performance electronic and spintronic applications. Incorporating narrow-band electronic correlation effects could further promote tunability, though mass renormalization may impact carrier mobility. It is therefore challenging to identify a vdW material with both high mobility and strong correlation. Herein, by a combination of optical spectroscopy and high-field quantum-oscillation measurements, we observe significant effective-mass enhancement in CeTe3 at low temperature, arising from not only the band-structure modulation by antiferromagnetic ordering but also the narrow-band correlation effect. Despite the mass enhancement, the quantum mobility surprisingly increases and reaches $\sim$2403 cm2/Vs, likely benefiting from topological protection. Remarkably, these unique properties are maintained in atomically thin nanoflakes with quantum mobility enhanced to $\sim$3158 cm2/Vs. Thus, CeTe3 emerges as a promising vdW antiferromagnetic metal with high-mobility heavy quasiparticles, potentially unlocking new device concepts.
@@ INTRODUCTION
Two-dimensional (2D) van der Waals (vdW) materials have been an active research area in condensed matter physics and materials science by virtue of their unique weak interlayer coupling, high carrier mobility, and strong tunability, hosting a rich variety of quantum phenomena such as 2D magnetism 2D-heterostructure ; 2D magnetic ; 2D: vdw-Nature , superconductivity SC-2D ; graphene-Nature , strong excitonic effects exciton-2D , topological states 2D-topo ; topological insulators , etc. These properties underpin promising applications in next-generation spintronic and electronic technologies. To further amplify the tunability of these systems, an effective strategy is to introduce strong electronic correlations - particularly by incorporating fluctuating $f$ electrons (e.g., Ce-$4f^{1}$, Yb-$4f^{13}$). The strong hybridization between localized $f$ electrons and itinerant conduction ($c$) electrons (i.e., $c$-$f$ Kondo hybridization) can give rise to a renormalized narrow-band electronic structure (Figure 1A), which is highly sensitive to tuning parameters such as temperature, doping, pressure, and magnetic field. This approach opens a new frontier: vdW-layered heavy-fermion systems CeSiI-Nature , which offer an unprecedented opportunity to integrate strong electron-electron interactions into a 2D framework, and bridge the fields of correlated electron physics and layered quantum materials, potentially unlocking new device concepts. However, as stronger correlations typically come at the cost of lower mobility due to an increased quasiparticle effective mass ($m^{*}$), following $\mu=e\tau/m^{*}$ (where $\tau$ is the relaxation time), therefore, it is extremely challenging to tailor a high-mobility vdW material within the context of strong correlation.
@@
(D) Temperature-dependent resistivity of bulk CeTe3 (pink) and LaTe3 (blue). The inset shows $\rho_{4f}$ ($\equiv\rho_{\text{Ce}}-\rho_{\text{La}}$), the resistivity contribution from scattering by $4f$ electrons. The dashed line signifies the $-\log T$ behavior due to incoherent Kondo scattering.
(E) Low-temperature resistivity and magnetic susceptibility both show the AFM1 transition at $T_{N1}=3.1$ K. The AFM2 transition at $T_{N2}$ is beyond the scope of our measurements.
The rare-earth tritelluride family $R$Te3 ($R$ = La-Tm, Y) has recently garnered extensive attention for coexistence of multiple charge-density-wave (CDW) orders RTe3-breakdown ; RTe3-OS ; RTe3-CDW ; CeTe3-STM ; CeTe3-CDW ; CeTe3-gap , pressure- and intercalation-induced superconductivity CeTe3-Pressure ; RTen-Pd inter , rich magnetic states Magnetic-TbTe3 ; Magnetic-RTe3 ; Magnetic-CeTe3 , and notably high mobility GdTe3-Sci adv ; LaTe3-interplay . Among them, CeTe3 is of particular interest in that the Ce-$4f^{1}$ configuration makes it a candidate 2D Kondo lattice. Yet, the role of electronic correlation in CeTe3 remains poorly understood. Furthermore, unlike LaTe3 and heavier lanthanide counterparts - where quantum-oscillation measurements have identified multiple well-resolved Fermi pockets - the Fermi surface (FS) topology of CeTe3 is still unclear CeTe3-QO1 ; CeTe3-QO2 ; CeTe3-QO3 ; RTe3-breakdown . This gap in knowledge obscures a full understanding of its quantum transport properties and the role of correlation effects.
In this work, we perform a comprehensive study of bulk CeTe3 via optical spectroscopy and high-field quantum oscillations. Besides the known low-frequency Fermi pocket ($\alpha$), we identify four additional pockets ($\beta$, $\gamma$, $\eta$, and $\delta$) with larger cross-sections, completing the bulk FS topology. A significant enhancement of carrier effective mass is observed around the onset of Kondo effect ($T_{K}^{on}\sim 17$ K), lending further support to the existence of moderately heavy fermions in this 2D Kondo-lattice compound. A second mass enhancement occurs at the antiferromagnetic (AFM) transition, accompanied by an unexpected increase in quantum mobility and a 0-to-$\pi$ Berry phase shift, suggesting a topological reconstruction of the Fermi surface. We further extend our investigation to the atomically thin limit. Strikingly, all the key features are retained in few-monolayer nanoflakes. These results establish CeTe3 as a rare vdW system that simultaneously hosts strong electronic correlation and high mobility even at the 2D limit, and thus lay a foundation for designing miniaturized spintronic devices leveraging strong electron correlations and robust magnetic order.
@@ RESULTS
CeTe3 crystallizes in an orthorhombic structure with the space group Cmcm (No. 63), as shown in Figure 1B. It consists of Ce-Te slabs sandwiched by Te-bilayer square nets stacking along the $\mathbf{b}$-axis (perpendicular to the Te nets). The adjacent Te-Te layers are weakly coupled by vdW force, and can be mechanically exfoliated into nanoflakes. The contributions to the FS are primarily from $p_{x}$ and $p_{z}$ orbitals of Te2 and Te3 in the 2D square nets, rendering a 2D characteristic (Figure 1C) CeTe3-CPB . The temperature-dependent resistivity of bulk CeTe3 is displayed in Figure 1D, where the results of its non-$4f$ reference, LaTe3, are also provided for comparison. In general, CeTe3 exhibits a good metallic behavior with a large residual resistivity ratio [RRR $\equiv\rho(300\text{K})/\rho(2\text{K})=40$], indicating the high quality of the single crystal. The scatterings contributed from $4f$ electrons, $\rho_{4f}$, can be separated out by subtracting the resistivity of LaTe3 (inset to Figure 1D). Around 100 K, a noticeable bump is observed in $\rho_{4f}(T)$, which should be ascribed to the scatterings between crystalline-electric-field split doublets, as is the case in most Ce-contained compounds LuoY-CeNi2As2 ; ChenK-CeCo2Ga8 . Below $\sim 17$ K, Kondo effect sets in due to the $c$-$f$ hybridization. Initially, the size of Kondo clouds (cf. Figure 1A) is small, so the Kondo scatterings are incoherent, characterized by the $-\log{T}$ law in $\rho_{4f}(T)$. Upon cooling, the Kondo coupling strengthens, and Kondo clouds communicate with each other, yielding the development of Kondo coherence that is manifested by a broad maximum in $\rho_{4f}(T)$ near $T_{coh}=5.5$ K. A clear inflection is seen at $T_{N1}=3.1$ K, due to the reduction of spin scattering in the AFM ordered state. The transition at $T_{N1}$ (denoted by AFM1) is also manifested in the magnetic susceptibility (Figure 1E). We are not able to see the second AFM transition at $T_{N2}=1.3$ K (AFM2) Magnetic-CeTe3 that is lower than the base temperature (1.5 K) of our measurements. It should be mentioned that previous specific heat measurements revealed a large Sommerfeld coefficient $\sim 400$ mJ/mol K2 in the normal state Magnetic-CeTe3 , suggestive of mass enhancement due to electronic correlation effect. All these confirm CeTe3 as a candidate vdW AFM dense Kondo lattice.
@@
(A) Drude-Lorentz fitting of $\sigma_{1}(\omega)$, taking $T=20$ K as an example. The three Lorentz components are depicted by dashed lines, whereas the two Drude components are marked in color-filled areas. The inset shows the temperature dependence of DC conductivity $\sigma_{dc}$ and the total contribution of Drude terms, $\sigma$-Drude, in the limit $\omega\rightarrow 0$.
@@
(B) Left, temperature dependence of $\omega_{p}^{2}\equiv\omega_{p1}^{2}+\omega_{p2}^{2}$; Right, the enhancement ratio of average effective mass of charge carriers, $\langle m^{*}\rangle/\langle m^{*}_{\text{300K}}\rangle$. The inset shows the sum of hole and electron carrier concentrations; data from Hall effect analysis, cf Figures S4-S5.
@@
To further look into the electronic correlation effect in CeTe3, we measured the optical spectroscopy at various temperatures [see Figure S2 in Supplemental Information (SI) for raw data]. In the real part of the optical conductivity, $\sigma_{1}(\omega)$, the features for the CDW gaps reported previously CeTe3-OS , e.g. the pronounced peak structure around 5000 cm-1 for all the measured temperatures and a second smaller peak centered near 2800 cm-1 present only for below 200 K, are well reproduced in our data. Here we mainly focus on the mass renormalization, and this can be obtained by the quantitative analysis of the optical conductivity. As shown in Figure 2A, $\sigma_{1}(\omega)$ can be well fitted to the Drude-Lorentz dispersion model Dressel-2002 ; Dai-La3Ni2O7 ,
$\sigma_{1}(\omega)=\frac{2\pi}{Z_{0}}\left[\sum_{k}\frac{\omega_{p,k}^{2}}{\tau_{k}(\omega^{2}+\tau_{k}^{-2})}+\sum_{i}\frac{\gamma_{i}\omega^{2}\Omega_{i}^{2}}{(\omega_{i}^{2}-\omega^{2})^{2}+\gamma_{i}^{2}\omega^{2}}\right],$ (1)
@@
where $Z_{0}=376.73~{}\Omega$ is vacuum impedance. The first term of Eq. (1) represents the sum of the Drude responses for free charge carriers (intraband transition), with $\omega_{p,k}$ and $1/\tau_{k}$ being the $k$-th component of plasma frequency and quasiparticle scattering rate, respectively. The second term of Eq. (1) is the sum of the Lorentz responses for localized electrons (interband transition) in which $\omega_{i}$, $\gamma_{i}$, and $\Omega_{i}$ stand for the resonance frequency, damping, and strength of the $i$-th oscillator, respectively. The fit employs two Drude terms, manifested respectively by the areas in green and blue in Figure 2A. The total Drude contribution in the limit $\omega\to 0$ matches well with the $\sigma_{dc}(T)$ profile, where $\sigma_{dc}\equiv 1/\rho$ is the DC conductivity (inset of Figure 2A). Other key parameters obtained from this analysis are $\omega_{p1}^{2}$ and $\omega_{p2}^{2}$, the sum of which (denoted as $\omega_{p}^{2}$) is displayed in Figure 2B as a function of temperature. One salient feature is that $\omega_{p}^{2}$ is essentially constant above 30 K, but decreases substantially below $T_{K}^{on}\sim 17$ K, manifesting an enhancement of quasiparticle effective mass due to Kondo hybridization. To see this more clearly, we estimate the average effective mass via $\langle m^{*}\rangle=\frac{Z_{0}ne^{2}}{2\pi\omega_{p}^{2}}$, where $n$ is the total carrier concentration that can be derived by Hall effect measurements [Figures S4 and S5 in SI]. A tiny decrease in $\langle m^{*}\rangle$ is observed from 300 K to 150 K (Figure 2B), likely due to the opening of the second CDW gap around 200 K CeTe3-OS . Below 100 K, $\langle m^{*}\rangle$ gradually increases, while a substantial enhancement is observed around $T_{K}^{on}$. At the lowest temperature of our optical measurements, 5 K, the enhancement ratio $\langle m^{*}\rangle/\langle m^{*}_{\text{300~{}K}}\rangle$ reaches $\sim 3.6$. Similar behavior has been reported in CeAlGe CeAlGe-OS ; WangZ-CeAlGe-NMR , a candidate of Kondo Weyl semimetal. This places CeTe3 in a regime with moderate electronic correlation. Though the factor of mass renormalization is moderate when compared with classic heavy-fermion metals, it is much larger than traditional gate-tuned 2D systems such as GaAs/AlGaAs heterostructure TanY-GaAsPRL2005 , bilayer graphene ZouK-BGraphenePRB2011 , etc. Note that earlier angle-resolved photoemission spectroscopy (ARPES) measurements on CeTe3 also revealed a weakly dispersive quasiparticle band near the Fermi level at low temperature CeTe3-CPB , in consistency with mass enhancement due to Kondo hybridization.
Quanum oscillations of bulk CeTe3
@@
Before running into our quantum oscillation results, it is necessary to put forward a qualitative sketch of Fermi-surface structure (Figure 1C). Due to the larger unit cell in the 3D crystal structure compared with the single square nets, the Brioullin zone (BZ) is reduced (from black dashed square to green dashed square). Correspondingly, the original FS (black solid lines) is “folded at the border of the reduced BZ. The nesting of the FS in the presence of CDW order not only removes FS elements but also creates new fractions CeTe3-FS ; NdTe3-FS , which eventually results in multiple FS pockets. Although multiple Fermi sheets were confirmed in other members of $R$Te3 by quantum oscillations RTe3-breakdown , no such results have been reported in CeTe3. Since enhanced effective mass typically means lower Fermi velocity, shorter transport mean free path and smaller carrier mobility, presumably, high magnetic field will be required to map out its FS topology by quantum oscillations.
@@
(D) FFT amplitudes as functions of $T$, and their LK fits (solid lines). Data of $\alpha$ are from SdH (4 - 16 T), while the remaining are from dHvA (20 - 54 T). An abrupt up-turn is clearly seen at $T_{N1}$, especially for the $\alpha$ pocket, which gives rise to different effective masses for above and below $T_{N1}$.
(E) Landau fan index diagram of the $\alpha$ oscillation at 5 K (red) and 1.5 K (blue). The solid lines represent the linear fittings that yield the intercepts $-0.448$ and $0.047$, respectively.

### v2

# Kondo-coupled van der Waals antiferromagnet with high-mobility quasiparticles

@@
Kondo-coupled van der Waals antiferromagnet with high-mobility quasiparticles
Hai Zeng1∗, Yang Zhang1∗, Bingke Ji2111These authors contribute equally to this work, Jiaqiang Cai3, Shuo Zou1, Zhuo Wang1, Chao Dong1, Kangjian Luo1, Yang Yuan1, Kai Wang4, Jinglei Zhang3, Chuanyin Xi3, Junfeng Wang1, Liang Li1,5, Yaomin Dai2222Electronic address: ymdai@nju.edu.cn, Jing Li1333Electronic address: jing_li@hust.edu.cn, and Yongkang Luo1,5,6444Electronic address: mpzslyk@gmail.com 1Wuhan National High Magnetic Field Center and School of Physics, Huazhong University of Science and Technology, Wuhan 430074, China; 2School of Physics, Nanjing University, Nanjing 210023, China; 3Anhui Key Laboratory of Low-Energy Quantum Materials and Devices, High Magnetic Field Laboratory, Hefei Institutes of Physical Science, Chinese Academy of Sciences, Hefei 230031, China; 4School of Physics and Electronics, Henan University, Kaifeng 475004, China; 5State Key Laboratory of Advanced Electromagnetic Technology, Huazhong University of Science and Technology, Wuhan 430074, China; 6Lead contact
(November 13, 2025)
SUMMARY: Two-dimensional van der Waals (vdW) materials exhibit high carrier mobility and tunability, making them suitable for low-power, high-performance electronic and spintronic applications. Incorporating narrow-band electronic correlation effects could further promote tunability, though mass renormalization may impact carrier mobility. It is therefore challenging to identify a vdW material with both high mobility and strong correlation. Herein, by a combination of optical spectroscopy and high-field quantum-oscillation measurements, we observe significant effective-mass enhancement in CeTe3 at low temperature, arising from not only the band-structure modulation by antiferromagnetic ordering but also the narrow-band correlation effect. Despite the mass enhancement, the quantum mobility surprisingly increases and reaches $\sim$2403 cm2/Vs, likely benefiting from topological protection. Remarkably, these unique properties are maintained in atomically thin nanoflakes with quantum mobility enhanced to $\sim$3158 cm2/Vs. Thus, CeTe3 emerges as a promising Kondo-coupled vdW antiferromagnetic metal with high-mobility quasiparticles, potentially unlocking new device concepts.
@@
Highlights
$\bullet$ vdW CeTe3 provides a rare example of strongly-correlated AFM metal with high mobility
$\bullet$ Complete Fermi surface topology of CeTe3 is mapped out by quantum oscillations
$\bullet$ Two mechanisms for quasiparticle-mass enhancement are confirmed in CeTe3
$\bullet$ High mobility, AFM order, and correlation maintained in atomically thin nanoflakes
@@ INTRODUCTION
Two-dimensional (2D) van der Waals (vdW) materials have been an active research area in condensed matter physics and materials science by virtue of their unique weak interlayer coupling, high carrier mobility ($\mu$), and strong tunability, exhibiting a rich variety of quantum phenomena, such as 2D magnetism burch_2018 ; mak_2019 , superconductivity saito_2017 ; torma_2022 ; pantaleon_2023 , strong exciton effects mak_2018 ; wang_2018 , and topological states culcer_2020 ; zhang_2022 , bearing a great prospect in high-efficient spintronic and electronic applications. These properties underpin promising applications in next-generation spintronic and electronic technologies. To further amplify the tunability of these systems, an effective strategy is to introduce strong electronic correlations - particularly by incorporating fluctuating $f$ electrons (e.g., Ce-$4f^{1}$, Yb-$4f^{13}$). The strong hybridization between localized $f$ electrons and itinerant conduction ($c$) electrons (i.e., $c$-$f$ Kondo hybridization) can give rise to a renormalized narrow-band electronic structure (Figure 1A), which is highly sensitive to tuning parameters such as temperature, doping, pressure, and magnetic field. This approach opens a new frontier: vdW-layered heavy-fermion systems CeSiI-Nature , which offer an unprecedented opportunity to integrate strong electron-electron interactions into a 2D framework, and bridge the fields of correlated electron physics and layered quantum materials, potentially unlocking new device concepts. However, as stronger correlations typically come at the cost of reduced mobility due to an increased quasiparticle effective mass ($m^{*}$), following $\mu=e\tau/m^{*}$ (where $\tau$ is the relaxation time), therefore, it is extremely challenging to tailor a high-mobility vdW material within the context of strong correlation.
@@
(D) Temperature-dependent resistivity of bulk CeTe3 (pink) and LaTe3 (black). The inset shows $\rho_{4f}$ ($\equiv\rho_{\text{Ce}}-\rho_{\text{La}}$), the resistivity contribution from scattering by $4f$ electrons. The dashed line signifies the $\log T$ behavior due to incoherent Kondo scattering.
(E) Low-temperature resistivity and magnetic susceptibility both show the AFM1 transition at $T_{\text{N1}}=3.1$ K. The AFM2 transition at $T_{\text{N2}}$ is beyond the scope of our measurements.
The rare-earth tritelluride family RRTe3 ($R$ = La-Tm, Y) has recently garnered extensive attention for coexistence of multiple charge-density-wave (CDW) orders RTe3-breakdown ; RTe3-OS ; RTe3-CDW ; CeTe3-STM ; CeTe3-CDW ; CeTe3-gap , pressure- and intercalation-induced superconductivity CeTe3-Pressure ; RTen-Pd inter , rich magnetic states Magnetic-TbTe3 ; Magnetic-RTe3 ; Magnetic-CeTe3 , and notably high mobility GdTe3-Sci adv ; LaTe3-interplay . Among them, CeTe3 is of particular interest in that the Ce-$4f^{1}$ configuration makes it a candidate 2D Kondo lattice. Yet, the role of electronic correlation in CeTe3 remains poorly understood. Furthermore, unlike LaTe3 and other lanthanide counterparts - where quantum-oscillation measurements have identified multiple well-resolved Fermi pockets - the Fermi surface (FS) topology of CeTe3 is less clear CeTe3-QO1 ; CeTe3-QO2 ; CeTe3-QO3 ; RTe3-breakdown ; Smith2024-CeTe3 . This gap in knowledge obscures a full understanding of its quantum transport properties and the role of correlation effects.
In this work, we perform a comprehensive study of bulk CeTe3 via optical spectroscopy and high-field quantum oscillations. Besides the known low-frequency Fermi pocket ($\alpha$), we identify four additional pockets ($\beta$, $\gamma$, $\eta$, and $\delta$) with larger cross-sections, completing the bulk FS topology. A significant enhancement of carrier effective mass is observed around the onset of Kondo effect ($T_{\text{K}}^{\text{on}}\sim 17$ K), lending further support to the existence of moderately heavy fermions in this 2D Kondo-lattice compound. A second mass enhancement occurs at the antiferromagnetic (AFM) transition, accompanied by an unexpected increase in quantum mobility and a 0-to-$\pi$ Berry phase shift, suggesting a topological reconstruction of the Fermi surface. We further extend our investigation to the atomically thin limit. Strikingly, all the key features are retained in few-monolayer nanoflakes. These results establish CeTe3 as a rare vdW system that hosts both strong electronic correlation and high charge mobility even at the 2D limit, and thus lay a foundation for designing miniaturized spintronic devices leveraging strong electronic correlation and robust magnetic order.
@@ RESULTS
CeTe3 crystallizes in an orthorhombic structure with the space group Cmcm (No. 63), as shown in Figure 1B. It consists of Ce-Te slabs sandwiched by Te-bilayer square nets stacking along the $\mathbf{b}$-axis (perpendicular to the Te nets). The adjacent Te-Te layers are weakly coupled by vdW force, and can be mechanically exfoliated into nanoflakes. The contributions to the FS are primarily from $p_{x}$ and $p_{z}$ orbitals of Te2 and Te3 in the 2D square nets, rendering a 2D characteristic (Figure 1C) CeTe3-CPB .
The temperature-dependent resistivity of bulk CeTe3 is displayed in Figure 1D, where the results of its non-$4f$ reference, LaTe3, are also provided for comparison. In general, CeTe3 exhibits a good metallic behavior with a large residual resistivity ratio [RRR $\equiv\rho(300\text{K})/\rho(2\text{K})=40$], indicating the high quality of the single crystal. The scatterings contributed from $4f$ electrons, $\rho_{4f}$, can be separated out by subtracting the resistivity of LaTe3 (inset to Figure 1D). Around 100 K, a noticeable bump is observed in $\rho_{4f}(T)$, which should be ascribed to the scatterings between crystalline-electric-field (CEF) split doublets, as is the case in most Ce-contained compounds LuoY-CeNi2As2 ; ChenK-CeCo2Ga8 . Below $\sim 17$ K, Kondo effect sets in due to the $c$-$f$ hybridization. Initially, the size of Kondo clouds (cf. Figure 1A) is small, so the Kondo scatterings are incoherent, characterized by the $\log{T}$ law in $\rho_{4f}(T)$. Upon cooling, the Kondo coupling strengthens, and Kondo clouds communicate with each other, yielding the development of Kondo coherence that is manifested by a broad maximum in $\rho_{4f}(T)$ near $T_{\text{coh}}=5.5$ K. A clear inflection is seen at $T_{\text{N1}}=3.1$ K, due to the reduction of spin scattering in the AFM ordered state CeTe3-QO3 . The transition at $T_{\text{N1}}$ (denoted by AFM1) is also manifested in the magnetic susceptibility (Figure 1E). We are not able to see the second AFM transition at $T_{\text{N2}}=1.3$ K (AFM2) Magnetic-CeTe3 that is lower than the base temperature (1.5 K) of our measurements. It should be mentioned that previous specific heat measurements revealed a large Sommerfeld coefficient $\sim 400$ mJ/mol K2 in the paramagnetic state Magnetic-CeTe3 , indicative of mass enhancement due to electronic correlation effect. All these suggest CeTe3 as a candidate vdW AFM dense Kondo lattice.
@@
(A) Drude-Lorentz fitting of $\sigma_{1}(\omega)$, taking $T=20$ K as an example. The three Lorentz components are depicted by dashed lines, whereas the two Drude components are marked in color-filled areas. The inset shows the temperature dependence of DC conductivity $\sigma_{\text{dc}}$ and the total contribution of Drude terms, $\sigma_{\text{Drude}}$, in the limit $\omega\rightarrow 0$.
@@
(B) Left, temperature dependence of $\omega_{\text{p}}^{2}\equiv\omega_{\text{p1}}^{2}+\omega_{\text{p2}}^{2}$; Right, temperature dependent average effective mass of charge carriers, $\langle m^{*}\rangle$. For comparison, the $\langle m^{*}\rangle$ of LaTe3 is also shown. The inset shows the mass-enhancement ratio, $\langle m^{*}_{\text{Ce}}\rangle/\langle m^{*}_{\text{La}}\rangle$.
@@
To further look into the electronic correlation effect in CeTe3, we measured the optical spectroscopy at various temperatures [see Figure S2 in Supplemental Information (SI) for raw data]. In the real part of the optical conductivity, $\sigma_{1}(\omega)$, the features for the CDW gaps reported previously sacchetti2006 ; sacchetti2007 ; CeTe3-OS , e.g. the pronounced peak structure around 5000 cm-1 for all the measured temperatures and a second smaller peak centered near 2800 cm-1 present only for below 200 K, are well reproduced in our data. Here we mainly focus on the mass renormalization, and this can be obtained by the quantitative analysis of the optical conductivity. As shown in Figure 2A, $\sigma_{1}(\omega)$ can be well fitted to the Drude-Lorentz dispersion model Dressel-2002 ; Dai-La3Ni2O7 ,
$\sigma_{1}(\omega)=\frac{2\pi}{Z_{0}}\left[\sum_{k}\frac{\omega_{\text{p},k}^{2}}{\tau_{k}(\omega^{2}+\tau_{k}^{-2})}+\sum_{i}\frac{\gamma_{i}\omega^{2}\Omega_{i}^{2}}{(\omega_{i}^{2}-\omega^{2})^{2}+\gamma_{i}^{2}\omega^{2}}\right],$ (1)
@@
where $Z_{0}=376.73~\Omega$ is vacuum impedance. The first term of Eq. (1) represents the sum of the Drude responses for free charge carriers (intraband transition), with $\omega_{\text{p},k}$ and $1/\tau_{k}$ being the $k$-th component of plasma frequency and quasiparticle scattering rate, respectively. The second term of Eq. (1) is the sum of the Lorentz responses for localized electrons (interband transition) in which $\omega_{i}$, $\gamma_{i}$, and $\Omega_{i}$ stand for the resonance frequency, damping, and strength of the $i$-th oscillator, respectively. The fit employs two Drude terms, manifested respectively by the areas in green and blue in Figure 2A. The total Drude contribution in the limit $\omega\to 0$ matches well with the $\sigma_{\text{dc}}(T)$ profile, where $\sigma_{\text{dc}}\equiv 1/\rho$ is the DC conductivity (inset of Figure 2A). Other key parameters obtained from this analysis are $\omega_{\text{p1}}^{2}$ and $\omega_{\text{p2}}^{2}$, the sum of which (denoted as $\omega_{\text{p}}^{2}$) is displayed in Figure 2B as a function of temperature. One salient feature is that $\omega_{\text{p}}^{2}$ is essentially constant above 30 K, but decreases substantially below $T_{\text{K}}^{\text{on}}\sim 17$ K, manifesting an enhancement of quasiparticle effective mass due to Kondo hybridization. To see this more clearly, we estimate the average effective mass via $\langle m^{*}\rangle=\frac{Z_{0}ne^{2}}{2\pi\omega_{\text{p}}^{2}}$, where $n$ is the total carrier concentration that can be derived by Hall effect measurements [Figures S5 and S6 in SI]. A tiny decrease in $\langle m^{*}\rangle$ is observed from 300 K to 150 K (Figure 2B), likely due to the opening of a second CDW gap around 200 K CeTe3-OS . Owing to the multiple CDW gaps, on the whole, $\langle m^{*}\rangle$ is small for all temperatures below 300 K. However, the remnant quasiparticles still exhibit prominent mass renormalization. Below 100 K, $\langle m^{*}\rangle$ gradually increases, while a substantial enhancement is observed around $T_{\text{K}}^{\text{on}}$. Recently, there has been accumulating evidence that fluctuating $c$-$f$ hybridization induced narrow bands start to appear near $T_{\text{K}}^{\text{on}}$ in prototypical Kondo-lattice compounds (e.g. CeCoIn5, CeCo2Ga8, etc), as exemplified by a series of angle-resolved photoemission spectroscopy (ARPES) ChenQY-PRB2017 ; KoitzschA-PRB2008 , ultrafast optical spectroscopy LiuY-PRL2020 , and planar Hall effect ZouS-CeCo2Ga8PHE2025 experiments. It, therefore, is reasonable to ascribe the $\langle m^{*}\rangle$ enhancement of CeTe3 around $T_{\text{K}}^{\text{on}}$ to the Kondo coupling effect. At the lowest temperature of our optical measurements, 5 K, the ratio $\langle m^{*}\rangle/\langle m^{*}_{\text{300K}}\rangle\approx 3.6$. In order to further demonstrate the effective mass enhancement, we also performed the same optical measurements on the non-$4f$ reference LaTe3, seeing Figure S4. The electronic correlation effect in CeTe3 can be directly visualized through the ratio $\langle m_{\text{Ce}}^{*}\rangle/\langle m_{\text{La}}^{*}\rangle$ that rises rapidly below $T_{\text{K}}^{\text{on}}$ and reaches a large value of $\sim 11.4$ at 5 K (inset to Figure 2B). Though the factors of $\langle m^{*}\rangle/\langle m^{*}_{\text{300K}}\rangle$ and $\langle m_{\text{Ce}}^{*}\rangle/\langle m_{\text{La}}^{*}\rangle$ are moderate when compared with classic heavy-fermion metals, they are much larger than traditional gate-tuned 2D systems such as GaAs/AlGaAs heterostructure TanY-GaAsPRL2005 , bilayer graphene ZouK-BGraphenePRB2011 , etc. All these place CeTe3 in a regime with moderate electronic correlation. Note that earlier ARPES measurements on CeTe3 also revealed a weakly dispersive quasiparticle band near the Fermi level at low temperature CeTe3-CPB , in consistency with the mass enhancement due to Kondo hybridization.
Quantum oscillations of bulk CeTe3
@@
Before running into our quantum oscillation results, it is necessary to put forward a qualitative sketch of Fermi-surface structure (Figure 1C). Due to the larger unit cell in the 3D crystal structure compared with the single square nets, the Brillouin zone (BZ) is reduced (from black dashed square to green dashed square). Correspondingly, the original FS (black solid lines) is “folded” at the border of the reduced BZ. The nesting of the FS in the presence of CDW order not only removes FS elements but also creates new fractions CeTe3-FS ; NdTe3-FS , which eventually results in multiple FS pockets. Although multiple Fermi sheets were confirmed by quantum oscillations in other members of RRTe3 RTe3-breakdown , no such results have been reported in CeTe3 CeTe3-QO1 ; CeTe3-QO2 ; CeTe3-QO3 ; Smith2024-CeTe3 . Since enhanced effective mass typically means lower Fermi velocity, shorter transport mean free path and smaller carrier mobility, presumably, high magnetic field will be required to map out its FS topology by quantum oscillations.
@@
(D) FFT amplitudes as functions of $T$, and their LK fits (solid lines). Data of $\alpha$ are from SdH (4 - 16 T), while the remaining are from dHvA (20 - 54 T). An abrupt upturn is clearly seen at $T_{\text{N1}}$, especially for the $\alpha$ pocket, which gives rise to different effective masses for above and below $T_{\text{N1}}$.
(E) Landau fan index diagram of the $\alpha$ oscillation at 5 K (red) and 1.5 K (blue). The solid lines represent the linear fittings that yield the intercepts 0.448-0.448 and $0.047$, respectively.